Pam is playing with red and black marbles. The number of red marbles she has is three more than twice the number of black marbles she has. She has 42 marbles in all how many red marbles does Pam have?
I got 29 red marbles but i am not sure if the equation i used to solve it was correct.
2x+3=42

Found 2 solutions by tushki, richwmiller:
Answer by tushki(2)

yes you did correctly

Answer by richwmiller(17219)

r+b=42
r=3+2b
Contrary to what my esteemed colleague has said, No, your equation was not correct.
You set up 3+2x=42
Your 3+2x should be the number of red marbles.

42 is the total number of marbles red + black
3+2b+b=42
3+3b=42
3b=39
b=13 black marbles
42-13=29 red marbles
I don't see how you calculated the correct 29 red marbles.